OpenStudy (anonymous):

A blue bike is $14 less than a red bike. The sum of their prices is $300. How much is the red bike?

OpenStudy (anonymous):

157

hero (hero):

Giving answers rather than explaining steps and ensuring understanding is not a good idea.

OpenStudy (anonymous):

r+r-14=300 r=157

OpenStudy (linna):

let's write price of a blue bike with x ... and price of red bike with y we have: x-14$=y x=y+14 and x+y=300$ substitute x into the second equation you get: (y+14)+y=300 2y=300-14 2y=286 y=143 now let's substitute: x+y=300 x+143=300 x=300-143 x=157 so price of the blue bike which we've written with X is 157 and price of the red bike which we've written with Y is 143 it also says that A blue bike is $14 less than a red bike so x-14=y 157-14=143 143=143 ---TRUE

hero (hero):

Proper Solution: If we let b = the price of the blue bike r = the price of the red bike Then, English statement: "A blue bike is $14 less than the red bike" becomes Mathematical Statement: b = r - 14 Also: English statement: "The sum of their prices is $300" becomes Mathematical statement b + r = 300 If we consider the mathematical statements only, then: b = r - 14 b + r = 300

OpenStudy (linna):

@Hero A blue bike is $14 less than a red bike. I think it should be b-14=r not b=r-14

hero (hero):

You should probably read the statement five more times before making a decision.

OpenStudy (linna):

ok .. I'll read it 5 more times :)

hero (hero):

Remember, the statement is: The blue bike is $14 less than the red bike.

hero (hero):

Hint: is means "equals"

OpenStudy (linna):

if blue bike costs 100$ and red bike 100$ b=r 100=100 100-14=100 86=100 still blue bike costs 14$ less than red bike !!

hero (hero):

Linna, you should be trying to convince yourself that the statement I wrote is true, not that your statement is true. I never disputed your statement.

OpenStudy (linna):

convince me.. give me an example. I want to know why I'm wrong. :)

hero (hero):

I never said you were wrong actually.

hero (hero):

But I will finish my solution just to convince you

OpenStudy (linna):

can't wait to see it !

hero (hero):

Proper Solution: If we let b = the price of the blue bike r = the price of the red bike Then, English statement: "A blue bike is $14 less than the red bike" becomes Mathematical Statement: b = r - 14 Also: English statement: "The sum of their prices is $300" becomes Mathematical statement b + r = 300 If we consider the mathematical statements only, then: b = r - 14 b + r = 300 Since we have two equations sharing the same variables, we can write both equations in terms of variable r: r = b + 14 r = 300 - b Now we can set r = r and solve for b: b + 14 = 300 - b 2b = 286 b = 143 Therefore the price of the blue bike is $143 This means that the price of the red bike must be $157 since: b = r - 14 143 = r - 14 143 + 14 = r 157 = r

OpenStudy (linna):

@Hero your solution seem to be correct, but... where I went wrong to get Blue bike 157 ??

hero (hero):

Your mathematical statement was incorrect to begin with.

hero (hero):

I have laid out the procedure of how to translate English Sentences to mathematical Statements

OpenStudy (linna):

anyway , can't say I wansn't close ! lol

OpenStudy (anonymous):

ok now leave that question, and move to another